Evaluate a) integral 0 to pi (dx/5-4 cos x) b) integral 0 to infinity (dx/(1+x^2)^3)

Evaluate

a) integral 0 to pi (dx/5-4 cos x)

b) integral 0 to infinity (dx/(1+x^2)^3)
Answer #1

(1) (93211 =1-2018) - 3-4+ asin Į dn - - - 1 Multiply by sec2 in humerator & denominator (sec?n=1 ttan ) It asin Me J sech daSelain = oman Saretynas + 2b(2-1) (9x2 +6771 de = secude s seczu du Salones = 3 sente + Fleet:) Sahh = (sechu de put n= tanuSo = tem 1x + Ź sh (2.taring) sin(2+er! (0) = p dhe = £ tarix + = So = ž (Panin + E) From (1) f ②, we get do = 3 (tan bot 1 4